# What is a Digital Twin Anyway? Deriving the Definition for the Built Environment from over 15,000 Scientific Publications

### v1

@@ Abstract
The concept of digital twins has attracted significant attention across various domains, particularly within the built environment. However, there is a sheer volume of definitions and the terminological consensus remains out of reach. The lack of a universally accepted definition leads to ambiguities in their conceptualization and implementation, and may cause miscommunication for both researchers and practitioners. We employed Natural Language Processing (NLP) techniques to systematically extract and analyze definitions of digital twins from a corpus of 15,000 full-text articles spanning diverse disciplines in the built environment. The study compares these findings with insights from an expert survey that included 52 experts. The study identifies concurrence on the components that comprise a “Digital Twin” from a practical perspective across various domains, contrasting them with those that do not, to identify deviations. We investigate the evolution of digital twin definitions over time and across different scales, including manufacturing, building, and urban/geospatial perspectives. We extracted the main components of Digital Twins using Text Frequency Analysis and N-gram analysis. Subsequently, we identified components that appeared in the literature and conducted a Chi-square test to assess the significance of each component in different domains. Our findings indicate that definitions differ based on the field of research in which they are conceived, but with many similarities across domains. One significant generalizable differentiation is related to whether a digital twin was used for High-Performance Real-Time (HPRT) or Long-Term Decision Support (LTDS) applications. We synthesized and contrasted the most representative definitions in each domain, culminating in a novel, data-driven definition specifically tailored for each context.
@@ 1 Introduction
The concept of a Digital Twin (DT) traces its origins to a book titled Mirrored Worlds published three decades ago (Gelernter, 1993). Various terms have been coined subsequently, e.g. Mirrored Space Model (MSM) and Information Mirroring Model (IMM), and later “Conceptual Ideal for Product Lifecycle Management (PLM)” by Grieves (2005), evolving the discourse with defining the main components of a DT. Since then, the term has evolved to encompass a broad spectrum of concepts and levels of fidelity, ranging from static models to dynamic simulations with real-time data integration and even autonomous capabilities (from a data visualization tool to a fully autonomous decision making system such as autonomous cars). For example, the term and concept of Digital Twin have been widely adopted in manufacturing industry (Leng et al., 2021; Jiang et al., 2021), and transcended it, reaching the built environment and urban/geospatial domains (Lehner and Dorffner, 2020; Dembski et al., 2020; Juarez et al., 2021; Sommer et al., 2023; Bartos and Kerkez, 2021; Chomiak-Orsa et al., 2023; White et al., 2021; Alva et al., 2022; Liu et al., 2023; Jeddoub et al., 2023; Metcalfe et al., 2024; Shirowzhan et al., 2020; Yan et al., 2019) including urban logistics and transportation (Belfadel et al., 2023; Kim and Bartos, 2024), urban-social interaction (Wang and Vu, 2023), participatory human-centric exploration (Lei et al., 2023c, 2024). conversion of Building Information Modeling (BIM), IoT, and GIS (Shi et al., 2023; Xia et al., 2022). Furthermore, according to more recent literature, the DT scope appears to have broadened to represent the entire lifecycle of a system, not just a single lifecycle phase (e.g., conceptual, design and planning, construction/manufacturing, operation and maintenance, and decomposition) (Boje et al., 2023) as well as a wide range of applications (Lei et al., 2023a). At the time of writing this paper (mid-2024), a query for the term ‘Digital Twin’ on ScienceDirect, a major database of peer-reviewed scholarly literature, yields 52,495 results. Although the majority of these results are dominated by Smart Manufacturing and Industrial Engineering (e.g., Industry 4.0), a total of 11,000 query results are related to the Built Environment (BE) [^1] We obtained this estimation by using the following boolean search: (‘Digital twin’ AND (‘Building’ OR ‘Built Environment’ OR ‘City’ OR ‘urban’ OR ‘smart building’ OR ‘BIM’ OR ‘architecture’) [^2] In the context of this paper, we refer to the BE as buildings, urban spaces/places and cities.. This multitude of literature demonstrates the intense and rapid interest in digital twin concept and technology and its applications in the built environment (Opoku et al., 2021; Pereira et al., 2021; Bolton et al., 2018; Xia et al., 2022). Despite the growing interest in DTs within the built environment, the concept has been reported to remain unclear for many practitioners and researchers (Osadcha et al., 2023; Chong et al., 2022), unlike in some other industries such as Manufacturing (Kritzinger et al., 2018) and Aerospace (Li et al., 2022) where it appears to be converging towards a well-established consensus.
@@
The rapid growth of interest in emergent technologies such as DTs can lead to fragmented knowledge about its core functionalities and key elements. This fragmentation often results in ambiguities across industry and research papers concerning its precise definition (Kritzinger et al., 2018; Bartsch et al., 2021; Cimino et al., 2019; Yang et al., 2021; Steindl et al., 2020), its essential components and minimum functionalities for a system to be considered a DT, and how it is applied specifically within the BE. For example, in discussions surrounding DTs, terms such as ‘data reference’, ‘software representation’, ‘living model’, ‘virtual replica’, and ‘visual asset management system’ are all used interchangeably. However, not all of them accurately capture the exact essence of a Digital Twin. Amplifying this confusion, researchers have defined DTs from one or more perspectives, often vaguely, leading to a range of interpretations (Table LABEL:table_definitions_review). Further, the ambiguity is exacerbated by its common use as a buzzword and the term is also affected by the expedient practice and trend of giving new names to existing concepts, simply to make them seem novel.
@@
Considering the inconsistent terminology and contrasting implementations, recent studies have introduced the concept of DT maturity stages (Klar et al., 2024; Haraguchi et al., 2024; Thaduri, 2024; Liu et al., 2024b), which is used to assess how advanced a DT is in terms of technology and functionality. These may range from basic and static geometric models to highly integrated systems with advanced features such as real-time capabilities and AI/ML functionalities supporting simulations. For example, Klar et al. (2024) and Evans et al. (2019) identified 5 maturity stages in City DTs: 1) Reality capture, 2) 2D maps or 3D models, 3) addition of real-time data, 4) two-way integration and interactions, and 5) autonomous operations and maintenance. Haraguchi et al. (2024) introduced the CITYSTEPS Maturity model which grades them according to eight maturity stages: 1) Preparation and planning stage; 2) 2D stage; 3) 3D static stage; 4) 3D dynamic stage; 5) Dynamically Integrated 3D Stage; 6) Real-time decision-making stage; 7) Autonomous decision-making stage; and 8) Real-time synchronization and autonomous implementation stage. Homa Masoumi and Pettit (2023) conducted a review on CDTs, highlighting their maturity levels and differentiating them from traditional 3D city models (Pang and Biljecki, 2022; Lei et al., 2023b; Paden et al., 2024), while identifying key research gaps, such as the need for advanced real-time data analytics and public participation in CDT development. These recent developments signal that the community is increasingly acknowledging that DTs may come in different shapes and flavours and considering Digital Twin as a broader umbrella term for a range of concepts and implementations of virtual replicas of the built environment. Nevertheless, an overarching definition and conceptualization of the term remain missing.
@@
In this paper, we conduct a massive systematic review coupled with an expert survey to shed light on the definition of a digital twin in the built environment. A systematic review methodology was used because it helps reduce bias in the selection and analysis of existing research, and it follows initiatives in other fields such as circular economy that have also grappled with similar terminological issues (Kirchherr et al., 2017; Reitsma, 2013). This review involves analyzing approximately one thousand Digital Twin mentions, interpretations, explanations, and definitions that have appeared in more than 15,000 publications from 2001 to January 2024. By systematically examining such a large sample size, we provided an empirical snapshot of current scholarly understandings of DTs, as well as the evolution of the definition over the years. We construct our approach on the premise that researchers are likely to choose definitions that align with their research focus or objectives, even if some of them do not align with the current predominant definitions. We then analyzed the definitions used in the Built Environment (Building, Urban domains), and we used Manufacturing domain as a benchmark for comparison to extract the DT components of each. The analysis aimed to understand the various perspectives on DTs, develop a definition based on the most common characteristics understood by researchers through practice and application, and ensure this definition was comprehensive enough to encompass a wide range of DT components, ultimately capturing the true essence of what constitutes a Digital Twin or in other words, of what BE researchers have been referring to as Digital Twins. We also consider the maturity levels of DT in different domains, recognizing that not all implementations require the same level of maturity. Finally, by deriving a definition, in our paper we question some instances in practice that gerrymander the definition of DT to masquerade their concepts as DT but significantly depart from the actual concept.
The approach to obtain clarity on terminology from a vast body of knowledge that we introduce in this paper is novel and our development, and a contribution that can be applied also in other domains.
@@ 2 Related work
As different domains recognize the potential of DTs and move toward implementations for required applications, within the paradigm of connected DTs, standardization plays an important role, especially as the concept of DTs continues to gain attention across different fields, particularly for its ability to continuously analyze incoming (i.e., real-time) data and perform data-driven simulations (Hananto et al., 2024; Lei et al., 2023a). Defining standards—which may relate to one or several elements of a DT and of its construction— is the scope of ongoing work (Wang et al., 2022a). Recently, the joint technical committee of the International Organization for Standardization (ISO) and International Electrotechnical Commission (IEC) published the standard ISO/IEC 30173:2023, Digital twin - Concepts and terminology which presents general descriptions of DT concepts, terminology and application overviews and aims to provide a common basis for understanding its concept and composition and be applicable in any and across domains (International Organization for Standardization, 2023). In the standard, a DT system is described as a hybrid entity or system of systems, with the DT elements illustrated within this context, which include the Target entity, which refers to something (physical or non-physical) having a distinct existence and is the subject of digital representation; the Digital entity, which is the computational entity comprised of data and procedural elements; and the Data connection between the two, and corresponding models, data and interfaces (e.g., APIs) involved—mirroring the DT concept model of Grieves (2014). Beyond these three, the standard further describes the other elements of the DT system as the Services, which refer to the capabilities (e.g. analysis, simulation, visualization, optimization, etc.) of the DT, the Infrastructure, which include the supporting components to ensure the DT’s operation (e.g., sensors, network, GIS, etc.) and the System aspects, which are concerned with the integration, governance, security, and privacy, and data ethics involved in the DT implementation. However, there has yet to be a DT reference model that unifies the ingredients necessary to build a DT (Raes et al., 2021). Thus, alongside unifying definitions, it is equally important to identify and unify the enabling components of DT development and implementation, especially toward goals of interoperability and scalability.
@@
While several academic reviews have attempted to identify or define DTs, these definitions are often aligned with specific research domains or adapted to fit particular applications. This multiplicity has resulted in a variety of definitions, some of which accurately represent the concept of a DT, while others do not and diverge from the mainstream. There is no single definition that fits all domains, and most existing definitions are meant to be universally applied. However, our research aims to redefine DTs on the basis of the researchers’ interpretation of the term within each specific domain, and we also challenge existing instances of (ostensible) DTs as we noticed papers that attempt to shoehorn certain instances into the DT concept. In doing so, we evaluate and establish definitions for each domain individually based on the essential components unique to each field and its users.
@@ 3 Methodology
Our approach leverages the credibility of existing literature, which has often undergone double-blind reviews by experts in various domains. This lends validity to several studies that qualify as a Digital Twin, even if they do not match standardized definitions. Therefore, we are deriving our definition from practice rather than theory, taking into account the extensive existing research (Figure 2). To validate our outcome, we conducted an expert survey from a panel of 52 experts in the BE DT domain. Additionally, we took into account the different maturity stages of DTs as defined by Klar et al. (2024) and Haraguchi et al. (2024).
@@
In this section, we first present our comprehensive approach to identifying the sample from various resources, including full-text publications and the expert survey. Then, we outline the methods used to extract and process approximately 800 definitions and interpretations of DTs from this sample. Our procedure includes a variety of Natural Language Processing (NLP) methods such as Regular Expressions (Regex) (Li et al., 2008), Sentence Embeddings (Le and Mikolov, 2014), Frequency Analysis and N-gram analysis (Cavnar et al., 1994), and Fuzzy Matching (Chaudhuri et al., 2003). Additionally, we employ Machine Learning approaches to group and cluster semantically similar definitions using k-means clustering and Large Language Models (LLMs). After that, we describe the methods used in extracting the DT components from different domains, namely, Building, Architecture, Urban, and Manufacturing using word cloud analysis and Large Language Models (LLMs). Finally, we test the significance of each component using Chi-square analysis. We maintained our methodology as systematic as possible to ensure it is generalizable to other research domains, which may also benefit from standardizing and analyzing terminology. Thus, the pipeline and code have been openly released on GitHub [^4] https://github.com/xxx/xxxx [will be added after publication]. Figures 3 and 4 summarize the methodology used in the current study.

### v2

@@ Abstract
The concept of digital twins has attracted significant attention across various domains, particularly within the built environment. However, there is a sheer volume of definitions and the terminological consensus remains out of reach. The lack of a universally accepted definition leads to ambiguities in their conceptualization and implementation, and may cause miscommunication for both researchers and practitioners.
We employed Natural Language Processing (NLP) techniques to systematically extract and analyze definitions of digital twins from a corpus of more than 15,000 full-text articles spanning diverse disciplines. The study compares these findings with insights from an expert survey that included 52 experts. The study identifies concurrence on the components that comprise a “Digital Twin” from a practical perspective across various domains, contrasting them with those that do not, to identify deviations. We investigate the evolution of digital twin definitions over time and across different scales, including manufacturing, building, and urban/geospatial perspectives. We extracted the main components of Digital Twins using Text Frequency Analysis and N-gram analysis. Subsequently, we identified components that appeared in the literature and conducted a Chi-square test to assess the significance of each component in different domains.
Our analysis identified key components of digital twins and revealed significant variations in definitions based on application domains, such as manufacturing, building, and urban contexts. The analysis of DT components reveal two major groups of DT types: High-Performance Real-Time (HPRT) DTs, and Long-Term Decision Support (LTDS) DTs. Contrary to common assumptions, we found that components such as simulation, AI/ML, real-time capabilities, and bi-directional data flow are not yet fully mature in the digital twins of the built environment. We derived two definitions for the Building/Architecture DT and the City/Urban DTs. Both definitions have a must-have components (such as spatial and temporal data updates) and good-to-have components such as prediction, AI, and Real-time data exchange. We also found that bi-directional flow (actuation) is not matured enough in the built environment.
One of the key findings is that the definition of digital twins has not yet reached its equilibrium phase, highlighting the need for ongoing revisions as technologies emerge or existing ones become obsolete. To address this, we introduce a novel, reproducible methodology that enables researchers to refine and adapt the current definitions in response to technological advancements or deprecations.
@@ 1 Introduction
The concept of a Digital Twin (DT) traces its origins to a book titled Mirrored Worlds published three decades ago (Gelernter, 1993). Various terms have been coined subsequently, e.g. Mirrored Space Model (MSM) and Information Mirroring Model (IMM), and later “Conceptual Ideal for Product Lifecycle Management (PLM)” by Grieves (2005), evolving the discourse with defining the main components of a DT. Since then, the term has evolved to encompass a broad spectrum of concepts and levels of fidelity, ranging from static models to dynamic simulations with real-time data integration and even autonomous capabilities (from a data visualization tool to a fully autonomous decision making system such as autonomous cars). For example, the term and concept of Digital Twin have been widely adopted in manufacturing industry (Leng et al., 2021; Jiang et al., 2021), and transcended it, reaching the built environment and urban/geospatial domains (Lehner and Dorffner, 2020; Dembski et al., 2020; Juarez et al., 2021; Sommer et al., 2023; Bartos and Kerkez, 2021; Chomiak-Orsa et al., 2023; White et al., 2021; Alva et al., 2022; Liu et al., 2023; Jeddoub et al., 2023; Metcalfe et al., 2024; Shirowzhan et al., 2020; Yan et al., 2019) including urban logistics and transportation (Belfadel et al., 2023; Kim and Bartos, 2024), urban-social interaction (Wang and Vu, 2023), participatory human-centric exploration (Lei et al., 2023c, 2024). conversion of Building Information Modeling (BIM), IoT, and GIS (Shi et al., 2023; Xia et al., 2022). Furthermore, according to more recent literature, the DT scope appears to have broadened to represent the entire lifecycle of a system, not just a single lifecycle phase (e.g., conceptual, design and planning, construction/manufacturing, operation and maintenance, and decomposition) (Boje et al., 2023) as well as a wide range of applications (Lei et al., 2023a). At the time of writing this paper (mid-2024), a query for the term ‘Digital Twin’ on ScienceDirect, a major database of peer-reviewed scholarly literature, yields 52,495 results. The majority of these results are dominated by Smart Manufacturing and Industrial Engineering (e.g., Industry 4.0). On a separate query, a total of 11,000 results are found to be related to the Built Environment (BE) [^1] We obtained this estimation by using the following boolean search: (‘Digital twin’ AND (‘Building’ OR ‘Built Environment’ OR ‘City’ OR ‘urban’ OR ‘smart building’ OR ‘BIM’ OR ‘architecture’) [^2] In the context of this paper, we refer to the BE as buildings, urban spaces/places and cities. which constitute different scales within the Architecture, Engineering, and Construction (AEC) domain. This multitude of literature demonstrates the intense and rapid interest in digital twin concept and technology and its applications in the built environment (Opoku et al., 2021; Pereira et al., 2021; Bolton et al., 2018; Xia et al., 2022). Despite the growing interest in DTs within the built environment, the concept has been reported to remain unclear for many practitioners and researchers (Osadcha et al., 2023; Chong et al., 2022), unlike in some other industries such as Manufacturing (Kritzinger et al., 2018) and Aerospace (Li et al., 2022) where it appears to be converging towards a well-established consensus.
@@
The rapid growth of interest in emergent technologies such as DTs can lead to fragmented knowledge about its core functionalities and key elements. This fragmentation often results in ambiguities across industry and research papers concerning its precise definition (Kritzinger et al., 2018; Bartsch et al., 2021; Cimino et al., 2019; Yang et al., 2021; Steindl et al., 2020), its essential components and minimum functionalities for a system to be considered a DT, and how it is applied specifically within the BE. For example, in discussions surrounding DTs, terms such as ‘data reference’, ‘software representation’, ‘living model’, ‘virtual replica’, and ‘visual asset management system’, ‘3D city models’ and ‘City Information Models (CIM)’ are all used interchangeably. However, not all of them accurately capture the exact essence of a Digital Twin. Amplifying this confusion, researchers have defined DTs from one or more perspectives, often vaguely, leading to a range of interpretations (Table LABEL:table_definitions_review). Further, the ambiguity is exacerbated by its common use as a buzzword and the term is also affected by the expedient practice and trend of giving new names to existing concepts, simply to make them seem novel.
@@
Considering the inconsistent terminology and contrasting implementations, recent studies have introduced the concept of DT maturity stages (Klar et al., 2024; Haraguchi et al., 2024; Thaduri, 2024; Liu et al., 2024b), which is used to assess how advanced a DT is in terms of technology and functionality. These may range from basic and static geometric models to highly integrated systems with advanced features such as real-time capabilities and AI/ML functionalities supporting simulations. For example, Klar et al. (2024) and Evans et al. (2019) identified 5 maturity stages in City DTs: 1) Reality capture, 2) 2D maps or 3D models, 3) addition of real-time data, 4) two-way integration and interactions, and 5) autonomous operations and maintenance. Haraguchi et al. (2024) introduced the CITYSTEPS Maturity model which grades them according to eight maturity stages: 1) Preparation and planning stage; 2) 2D stage; 3) 3D static stage; 4) 3D dynamic stage; 5) Dynamically Integrated 3D Stage; 6) Real-time decision-making stage; 7) Autonomous decision-making stage; and 8) Real-time synchronization and autonomous implementation stage. Homa Masoumi and Pettit (2023) conducted a review on CDTs, highlighting their maturity levels and differentiating them from traditional 3D city models (Pang and Biljecki, 2022; Lei et al., 2023b; Paden et al., 2024), while identifying key research gaps, such as the need for advanced real-time data analytics and public participation in CDT development. These recent developments signal that the community is increasingly acknowledging that DTs may come in different shapes and flavours and considering Digital Twin as a broader umbrella term for a range of concepts and implementations of virtual replicas of the built environment. Nevertheless, an overarching definition and conceptualization of the term in the built environment remain missing (Delgado and Oyedele, 2021; Davila Delgado and Oyedele, 2021; Jones et al., 2020) despite the substantial efforts invested to address this gap (Ellul et al., 2022; Stoter et al., 2021; Ellul et al., 2024; Callcut et al., 2021; Alva et al., 2022).
@@
In this paper, we conduct a large scale systematic review coupled with an expert survey to shed light on the definition of a digital twin in the built environment. A systematic review methodology was used because it helps reduce bias in the selection and analysis of existing research, and it follows initiatives in other fields such as circular economy that have also grappled with similar terminological issues (Kirchherr et al., 2017; Reitsma, 2013). This review involves analyzing approximately one thousand Digital Twin mentions, interpretations, explanations, and definitions that have appeared in more than 15,000 publications from 2001 to January 2024. By systematically examining such a large sample size, we provided an empirical snapshot of current scholarly understandings of DTs, as well as the evolution of the definition over the years. We construct our approach on the premise that researchers are likely to choose definitions that align with their research focus or objectives, even if some of them do not align with the current predominant definitions. We then analyzed the definitions used in the Built Environment (Building, Urban domains), and we used Manufacturing domain as a benchmark for comparison to extract the DT components of each. The analysis aimed to understand the various perspectives on DTs, develop a definition based on the most common characteristics understood by researchers through practice and application, and ensure this definition was comprehensive enough to encompass a wide range of DT components, ultimately capturing the true essence of what constitutes a Digital Twin or in other words, of what BE researchers have been referring to as Digital Twins. We also consider the maturity levels of DT in different domains, recognizing that not all implementations require the same level of maturity. Finally, by deriving a definition, in our paper we question some instances in practice that manipulate the definition of DT to align their concepts as DT but significantly depart from the actual concept.
The approach to obtain clarity on terminology from a vast body of knowledge that we introduce in this paper is novel and our development, and a contribution that can be applied also in other domains. The rest of the article is structured as follows. Section 2 provides a review of the literature on Digital Twins, highlighting the various definitions that have been proposed in the literature. Section 3 outlines the methodology used in this research, including data sources, data processing, and the methodology to derive a new definition of Digital Twin. Section 4 presents the results of the study, including the analysis of the components and definitions extracted from the literature and the expert survey. Section 5 discusses the findings, implications, and futures of the study, and Section 6 concludes the paper.
@@ 2 Related work
The concept of DTs continues to gain attention across different fields, particularly for its ability to continuously analyze incoming (i.e., real-time) data and perform data-driven simulations (Hananto et al., 2024; Lei et al., 2023a). As different domains recognize the potential of DTs and move toward implementations for various applications, standards play an important role in the creation of DTs, especially concerning data from different sources for use in various applications Ellul et al. (2024). For complex systems such as cities, the practical approach of digital twinning would be to begin with priority areas and scale over time which would require common frameworks for effective communication Raes et al. (2021).
@@
Defining standards—which may relate to one or several elements of a DT and of its construction— is the scope of ongoing work (Wang et al., 2022a). Recently, the joint technical committee of the International Organization for Standardization (ISO) and International Electrotechnical Commission (IEC) published the standard ISO/IEC 30173:2023, Digital twin - Concepts and terminology which presents general descriptions of DT concepts, terminology and application overviews and aims to provide a common basis for understanding its concept and composition and be applicable in any and across domains (International Organization for Standardization, 2023). In the standard, a DT system is described as a hybrid entity or system of systems, with the DT elements illustrated within this context, which include the Target entity, which refers to something (physical or non-physical) having a distinct existence and is the subject of digital representation; the Digital entity, which is the computational entity comprised of data and procedural elements; and the Data connection between the two, and corresponding models, data and interfaces (e.g., APIs) involved—mirroring the DT concept model of Grieves (2014). Beyond these three, the standard further describes the other elements of the DT system as the Services, which refer to the capabilities (e.g. analysis, simulation, visualization, optimization, etc.) of the DT, the Infrastructure, which include the supporting components to ensure the DT’s operation (e.g., sensors, network, GIS, etc.) and the System aspects, which are concerned with the integration, governance, security, and privacy, and data ethics involved in the DT implementation. However, there has yet to be a DT reference model that unifies the ingredients necessary to build a DT (Raes et al., 2021). Thus, alongside unifying definitions, it is equally important to identify and unify the enabling components of DT development and implementation, especially toward goals of interoperability and scalability.
@@
While several academic reviews have attempted to identify or define DTs, these definitions are often aligned with specific research domains or adapted to fit particular applications. This multiplicity has resulted in a variety of definitions, some of which accurately represent the concept of a DT, while others do not and diverge from the mainstream. There is no single definition that fits all domains, and most existing definitions are meant to be universally applied. However, our research aims to redefine DTs on the basis of the researchers’ interpretation of the term within each specific domain, and we also challenge existing instances of (ostensible) DTs as we noticed papers that attempt to forcibly fit certain instances into the DT concept. In doing so, we evaluate and establish definitions for each domain individually based on the essential components unique to each field and its users.
@@ 3 Methodology
Our approach leverages the credibility of existing literature, which has often undergone double-blind reviews by experts in various domains. This lends validity to several studies that qualify as a Digital Twin, even if they do not match standardized definitions. Accordingly, our methodology is grounded in both empirical observations and the extensive body of existing literature, combining theoretical insights with practical validation (Figure 2). To validate our outcome, we conducted an expert survey from a panel of 52 experts in the BE DT domain. Additionally, we considered different maturity stages of DTs as defined by Klar et al. (2024) and Haraguchi et al. (2024).
@@
In this section, we first present our approach to identifying the sample from various resources, including full-text publications and the expert survey. Then, we outline the methods used to extract and process approximately 800 definitions and interpretations of DTs from this sample. Our procedure includes a variety of NLP methods such as Regular Expressions (Regex) (Li et al., 2008), Sentence Embeddings (Le and Mikolov, 2014), Frequency Analysis and N-gram analysis (Cavnar et al., 1994), and Fuzzy Matching (Chaudhuri et al., 2003). Additionally, we employ Machine Learning approaches to group and cluster semantically similar definitions using k-means clustering and Large Language Models (LLMs). After that, we describe the methods used in extracting the DT components from different scales within the Architecture, Engineering, and Construction domain; namely, Building, Architecture, and Urban scales; in addition to Manufacturing domain as a dominant domain in this area. We used word cloud analysis and LLMs for that purpose. Then we test the significance of each component using Chi-square analysis. We maintained our methodology as systematic as possible to ensure it is generalizable to other research domains, which may also benefit from standardizing and analyzing terminology. Thus, the pipeline and code have been openly released on GitHub [^4] https://github.com/ualsg/dt-definitions. Figures 3 and 4 summarize the methodology used in the current study.